# UNEM: UNrolled Generalized EM for Transductive Few-Shot Learning

### v1

@@ D.2 Illustration and analysis of the learned hyperparameters
Illustration of the learned hyper-parameters
The evolutions of the learned hyper-parameters $\lambda^{(\ell)}$ and $T^{(\ell)}$ with respect to the layer index are illustrated in Figures 4 and 5 for some downstream image classification tasks. While Figure 4 shows that the learned hyper-parameters with CUB (ResNet18), mini-ImageNet (ResNet18), and mini-ImageNet (WRN28-10) have similar amplitudes, much different orders-of-magnitude are observed with vision-language models as shown in Figure 5 for some test datasets. Let us recall that the different learned hyper-parameters, for all datasets, are available at https://anonymous.4open.science/r/UNEM.
Analysis of the learned hyper-parameters

### v2

@@ E Illustration and analysis of the learned hyperparameters
$\bullet$ Illustration of the learned hyper-parameters
The evolutions of the learned hyper-parameters $\lambda^{(\ell)}$ and $T^{(\ell)}$ with respect to the layer index are illustrated in Figures 4 and 5 for some downstream image classification tasks. While Figure 4 shows that the learned hyper-parameters with CUB (ResNet18), mini-ImageNet (ResNet18), and mini-ImageNet (WRN28-10) have similar amplitudes, much different orders-of-magnitude are observed with vision-language models as shown in Figure 5 for some test datasets. Let us recall that the different learned hyper-parameters, for all datasets, are available at https://github.com/ZhouLong0/UNEM-Transductive.
$\bullet$ Analysis of the learned hyper-parameters